# PARCv2: Physics-aware Recurrent Convolutional Neural Networks for Spatiotemporal Dynamics Modeling

### v2

@@ 1 Introduction
Physics-aware deep learning (PADL) has been gaining attention for the simulation of spatiotemporal dynamics, i.e., time-evolving fields of physical systems. The physical behaviors of continua are represented as a system of partial differential equations (PDE), expressing the rate of change $\frac{\partial\mathbf{x}}{\partial t}$ of a field $\mathbf{x}(t,\mathbf{r})$ as a function $f(t,\mathbf{x},\mathbf{y},\nabla\mathbf{x},\nabla\mathbf{y},\cdots)$ where $t$ is time, $\mathbf{r}\in d$ is a position vector, $\mathbf{x}$ is the evolving field variable, and $\mathbf{y}(t,\mathbf{r})$ is a covariate field that influences the evolution of $\mathbf{x}$. As a simple case, the heat equation $\frac{\partial\mathbf{x}}{\partial t}=\nabla^{2}\mathbf{x}$ is one such governing PDE that describes the diffusion of heat and appears frequently in various computer vision and machine learning literature.
@@
As the complexity of the physics displayed by a system increases, the governing PDEs correspondingly increase in number and complexity, with nonlinear coupling between the dependent variables and source terms inducing a wide range of spatial and temporal scales. The constitutive laws describing the material response can also become complex nonlinear differential equations. Furthermore, depending on the physics problem, governing PDEs relating $\mathbf{x}$ and $\mathbf{y}$ may not even be available in an explicit form, and hence the equations governing the evolution of these variables may need to be learned from data.
@@
Interesting machine-learning challenges may emerge when there are large, localized spatial gradients in $\mathbf{x}$ and $\mathbf{y}$. Examples may include shocks, fast transients, collapsing interfaces, reaction fronts, and evolving internal/external boundaries interacting with the evolving field. An archetype of such problems is the shock-induced initiation of porous energetic materials (Perry et al., 2018; Mang and Hjelm, 2021; Mi et al., 2020), as in the benchmark problem presented in Section 4.3. When exposed to external shock loading, pores in these materials collapse and release strong localized energy, forming high-temperature regions called “hotspots.” In this process, both the traveling shock and the energy localization near hotspots render intense spatiotemporal gradients in the evolving temperature and pressure fields. Collapsing pore boundaries also interact with these evolving fields, producing intricate sharp patterns on the reaction front. Hence, accurately predicting such sharp and complex features evolving over time is critical to capture the physics of ignition and reaction in shocked reactive materials.
@@
Simulating the intricate physics of systems with fast transients and evolving spatiotemporal fields demands high-fidelity numerical techniques, incurring substantial computational costs and resulting in limited training data for machine learning. To address this challenge, previous works have used the squared sum of residuals of PDEs (known as physics-informed loss) to regularize the training. However, strong regularization can introduce undesired "smoothing effects," generate conflicting network gradients during training, and may be inapplicable when direct mathematical relationships between the evolving field $\mathbf{x}$ and covariate fields $\mathbf{y}$ are unavailable. Given these practical scenarios, our goal is to model the governing PDEs of spatiotemporal dynamic systems using neural networks. This enables us to either approximate the solution of a complex PDE system (Sections 4.1 &\&& 4.2) or discover unknown governing PDEs from data (Section 4.3).
@@ 2 Related Works
Spectral bias is another current challenge, arising from the tendency of PINN models to favor low-frequency modes in field evolution problems. This is largely due to unbalanced sampling of high-frequency patterns during the evaluation of PDE residuals, and can thus be somewhat mitigated by incorporating adaptive weighting (Wang et al., 2022) or Fourier/spectral features (Wang et al., 2021). However, for physics problems with high Kolmogorov n-width, e.g., ones exhibiting high advection dominance, spatial localization phenomena, discontinuous phases, and/or sharp gradients, learning bias models still fail to learn crucial intricate details (Mojgani et al., 2023).
@@
In contrast to enforcing physical constraints by minimizing PDE residuals, inductive bias approaches attempt to embed prior physics knowledge within the network’s computational graph itself. For example, equivariant neural networks (Weiler et al., 2018; Batzner et al., 2022; Gong et al., 2023; Ma et al., 2023) preserve the symmetry under a group operation (e.g., the standard Euclidean group) by redefining the convolution filters in such a way that the convolution responses are equivariant under certain geometric transformations.
However, little to no previous work has been done for inductive bias modeling of more general nonlinear field evolution problems that extend beyond problems exhibiting group symmetry. To the best of our knowledge, the physics-aware recurrent convolutional neural networks, or PARC (Nguyen et al., 2023a, b) is the only method in this category. PARC aims to model the governing equation of a simplified dynamical system $\frac{\partial\mathbf{x}}{\partial t}=f(t,\mathbf{x},\mathbf{y},\nabla\mathbf{x},\nabla\mathbf{y},\cdots)$ and its time integration using recursively connected convolutional neural network (CNN) blocks, called differentiator and integrator. The differentiator CNN approximates the function $f$ in the above expression. The time derivative $\frac{\partial\mathbf{x}}{\partial t}$ predicted by the differentiator CNN is then integrated by the integrator CNN, which approximates the integral operator using a CNN. Applied recursively, the differentiator-integrator architecture emulates the process by which traditional physics solvers obtain the solution of governing PDEs.
Despite certain successes, the previous work of Nguyen et al. (2023a) does not fully extend to a broader spectrum of PDEs, especially the ones in the regime of strong advection dominance. Nguyen et al. circumvented such a limitation by normalizing the velocity field and removing advection, but it is questionable if such a normalization strategy would generalize to more complex, transient physics phenomena.
@@ 3 Method
Here, we consider a wide group of advection-diffusion-reaction processes, in which the temporal evolution of a state field $\mathbf{x}$ in a continuum domain is the combination of advection $(\mathbf{u}\cdot\nabla)\mathbf{x}$ driven by a physical force represented by a (velocity) vector field $\mathbf{u}$, diffusion $(\nabla\cdot k\nabla)\mathbf{x}$ with diffusion coefficient $k$, and the generation/consumption of a source/sink (reaction) $\mathbf{R}(\mathbf{x},\mathbf{c})$. By explicitly incorporating the spatial gradient $\nabla\mathbf{x}$ of the state field and its divergence $\nabla\cdot\nabla\mathbf{x}$ into the differentiator architecture, we inform PARCv2 with physics priors (Section 3.1). Furthermore, a hybrid integration approach is introduced to leverage the benefits of numerical integration with data-driven integration (Section 3.2), enabling accurate and stable predictions over a long time period. Finally, we also present a two-stage training scheme that separates the process of learning the PDEs and the time integration (Section 3.3).
@@ 3.1 Architecture design
The dynamics of an advection-diffusion-reaction system can be described by the following PDE:
@@ 3.2 Hybrid integration
For physical systems with unsteady and fast transient characteristics, simulations can quickly deviate from the true dynamic trajectories as numerical and modeling errors accumulate during time integration. Traditional numerical methods for initial value problems such as the Runge-Kutta methods or the Adams methods take multiple steps in time to correct the error accumulation heuristically. However, unsteady and fast transient problems require a higher-order correction with a small time step, making it less tractable for our purpose of rapidly estimating the dynamics. The original PARC Nguyen et al. (2023a) and other works in literature  (Shen et al., 2020) explored the use of a deep neural network-based quadrature for time integration. However, as we will demonstrate in Sections 4.2 & 4.3, these purely data-driven integrators fail to predict advection features with sharp boundaries precisely.
@@
3.3 Learning the governing PDEs and the data-driven solver
@@ 4.1 Burgers’ equation
[…]
@@
4.2 Navier-Stokes equation for fluid flow around an obstacle
@@
Figure 4 displays the fluid velocity fields predicted by the models. Visually, all methods appear to capture the oscillatory vorticity patterns to a certain extent. However, the oscillatory phase of the vorticity patterns was most accurate in PARCv2, highlighting its stable long-time prediction capabilities. This observation is further supported by Table 1, where the RMSE of PARCv2 was 0.0727 $(m/s)$—roughly 1/4 of the values obtained by the other methods.
Interestingly, the original PARC, only with a purely data-driven integrator (fourth row in Figure 4), exhibited highly noisy predictions. This may be indicative of the rapid accumulation of network prediction errors in unsteady flows with sharp features. It is also worthwhile to compare PARCv2 (second row in Figure 4) with the original PARC with a numerical integrator (third row in the same figure). Compared to the purely data-driven integration case, the original PARC exhibited a more stable prediction with the numerical integrator. However, the original PARC still appears to be off from the ground truth in terms of the oscillatory phase of the vorticity pattern, while PARCv2 aligns well with the true phase. This result highlights the importance of the data-driven higher-order correction in the integrator. Furthermore, not only the phase but also the geometric pattern slightly deviates from the ground truth in the case of the original PARC, in contrast to the PARCv2 result where the geometric patterns are faithfully reproduced. This is indicative of the role of the spatial gradient terms newly introduced in the PARCv2 architecture.
@@ 5 Conclusion
These advantages underscore the significance and potential of the inductive bias research direction, positioning it as a promising complement to current mainstream learning bias approaches. It is, in fact, important to note that PARCv2 is not positioned as an exclusive alternative to learning-bias methods, as the incorporation of physics-informed loss during PARC training is still possible. Future research directions may explore the synergistic leverage of soft constraints imposed by the physics-informed loss and the inductive bias provided by the PARC differentiator-integrator architecture.
@@
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Burgers’ equation
Burgers’ equation is a convection-diffusion equation that occurs in a variety of applications such as fluid mechanics, acoustics, plasma physics, and traffic flows. With the presence of viscous (diffusion) term, the shock discontinuities are smoothed out and one can expect to obtain a well-behaved and smooth solution which will make the learning task less challenging for PARCv2 and other models. Therefore, learning to solve Burgers’ equation can be considered the most straightforward compared to the other two validation studies and was used as the first experiment for our validation.
@@
The fluid parameters, i.e., the Reynolds number $R$ and the initial velocity distribution $a$ and $w$ vary according to the values listed in Table 3. Note that the fluid parameters in the train-test split are designed to test the models’ abilities to not only interpolate within the training distributions but also extrapolate/generalize to conditions outside of the training distributions.
@@
The training data was obtained by solving Eq. 9 with numerical simulation using implicit backward Euler integration. Data was generated using the open-source code provided by [ref: LLNL LaSDI]. The original numerical simulation data includes 1500 snapshots and was down-sampled to 100 snapshots to be used for training. The spatial resolution of the numerical simulation data was kept during the training.

### v3

@@ 1 Introduction
Physics-aware deep learning (PADL) has been gaining attention for the simulation of spatiotemporal dynamics, i.e., time-evolving fields of physical systems. The physical behaviors of continua are represented as a system of partial differential equations (PDE), expressing the rate of change $\frac{\partial\mathbf{x}}{\partial t}$ of a field $\mathbf{x}(t,\mathbf{r})$ as a function $f(t,\mathbf{x},\mathbf{y},\nabla\mathbf{x},\nabla\mathbf{y},\cdots)$ where $t$ is time, $\mathbf{r}\in\mathbb{R}^{d}$ is a position vector, $\mathbf{x}$ is the evolving field variable, and $\mathbf{y}(t,\mathbf{r})$ is a covariate field that influences the evolution of $\mathbf{x}$. As a simple case, the heat equation $\frac{\partial\mathbf{x}}{\partial t}=\nabla^{2}\mathbf{x}$ is one such governing PDE that describes the diffusion of heat and appears frequently in various computer vision and machine learning literature.
@@
As the complexity of the physics displayed by a system increases, the governing PDEs increase in number and complexity accordingly, with nonlinear coupling between the dependent variables and source terms inducing a wide range of spatial and temporal scales. The constitutive laws describing the material response can also become complex nonlinear differential equations. Furthermore, depending on the physics problem, governing PDEs relating $\mathbf{x}$ and $\mathbf{y}$ may not even be available in an explicit form, and hence the equations governing the evolution of these variables may need to be learned from data.
@@
Particularly, interesting machine-learning challenges may emerge when there are large, localized spatial gradients in $\mathbf{x}$ and $\mathbf{y}$. Examples may include shocks, fast transients, collapsing interfaces, reaction fronts, and evolving internal/external boundaries interacting with the evolving field. An archetype of such problems is the shock-induced initiation of porous energetic materials (Perry et al., 2018; Mang and Hjelm, 2021; Mi et al., 2020), as in the benchmark problem presented in Section 4.3. When exposed to external shock loading, pores in these materials collapse and release strong localized energy, forming high-temperature regions called “hotspots.” In this process, both the traveling shock and the energy localization near hotspots render intense spatiotemporal gradients in the evolving temperature and pressure fields. Collapsing pore boundaries also interact with these evolving fields, producing intricate sharp patterns on the reaction front. Hence, accurately predicting such sharp and complex features evolving over time is critical to capture the physics of ignition and reaction in shocked reactive materials.
@@
Simulating the intricate physics of systems with fast transients and evolving spatiotemporal fields demands high-fidelity numerical techniques, incurring substantial computational costs and resulting in limited training data for machine learning. To address this challenge, previous works have used the squared sum of residuals of PDEs, known as physics-informed loss, to regularize the training. However, strong regularization can introduce undesired “smoothing effects,” generate conflicting network gradients during training, and may be inapplicable when direct mathematical relationships between the evolving field $\mathbf{x}$ and covariate fields $\mathbf{y}$ are unavailable. Given these practical scenarios, our goal is to model the governing PDEs of spatiotemporal dynamic systems using neural networks. This enables us to either approximate the solution of a complex PDE system (Sections 4.1 &\&& 4.2) or discover unknown governing PDEs from data (Section 4.3).
@@ 2 Related Works
Spectral bias is another current challenge, arising from the tendency of PINN models to favor low-frequency modes in field evolution problems. This is largely due to unbalanced sampling of high-frequency patterns during the evaluation of PDE residuals, and can thus be somewhat mitigated by incorporating adaptive weighting (Wang et al., 2022) or Fourier/spectral features (Wang et al., 2021). However, for physics problems with large Kolmogorov $n$-width, e.g., ones exhibiting high advection dominance, spatial localization phenomena, discontinuous phases, and/or sharp gradients, learning bias models still fail to learn crucial intricate details (Mojgani et al., 2023).
@@
In contrast to enforcing physical constraints by minimizing PDE residuals, inductive bias approaches attempt to embed prior physics knowledge within the network’s computational graph itself. For example, equivariant neural networks (Weiler et al., 2018; Batzner et al., 2022; Gong et al., 2023; Ma et al., 2023) preserve the symmetry under a group operation (e.g., the standard Euclidean group $SE(n)$) by redefining the convolution filters in such a way that the convolution responses are equivariant under certain geometric transformations.
However, little to no previous work has been done for inductive bias modeling of more general nonlinear field evolution problems that extend beyond problems exhibiting group symmetry. Finite volume neural network (FINN), introduced by Karlbauer et al. (2022), is among these minorities. In FINN, multiple neural network modules are used to model specific parts of the target PDE. These modules interact with each other in a distributed and compositional manner, aiming to model the physics governed by the target PDE. FINN demonstrated success in dealing with various types of PDEs, including Burgers, diffusion-reaction, and Allen–Cahn equations. However, one of the limitations of FINN is that the PDE constants, e.g., the diffusion coefficient, are treated directly as network parameters (either trainable or untrainable). This restricts the model’s expressiveness in modeling nonlinearities in PDEs and may limit its performance when dealing with PDEs with large variations of constants.
The physics-aware recurrent convolutional neural networks, or PARC (Nguyen et al., 2023a, b) is the another notable work in this category. PARC aims to model the governing equation of a simplified dynamical system $\frac{\partial\mathbf{x}}{\partial t}=f(t,\mathbf{x},\mathbf{y},\nabla\mathbf{x},\nabla\mathbf{y},\cdots)$ and its time integration using recursively connected convolutional neural network (CNN) blocks, called differentiator and integrator. The differentiator CNN approximates the function $f$ in the above expression. The time derivative $\frac{\partial\mathbf{x}}{\partial t}$ predicted by the differentiator CNN is then integrated by the integrator CNN, which approximates the integral operator using a CNN. Applied recursively, the differentiator-integrator architecture emulates the process by which traditional physics solvers obtain the solution of governing PDEs.
However, despite certain success, the previous work of Nguyen et al. (2023a) does not fully extend to a broader spectrum of PDEs, especially the ones in the regime of strong advection dominance. Nguyen et al. circumvented such a limitation by normalizing the velocity field and removing advection, but it is questionable if such a normalization strategy would generalize to more complex, transient physics phenomena.
@@ 3 Method
Here, we consider a wide group of advection-diffusion-reaction processes, in which the temporal evolution of a state field $\mathbf{x}$ in a continuum domain is the combination of advection $(\mathbf{u}\cdot\nabla)\mathbf{x}$ driven by a physical force represented with a (velocity) vector field $\mathbf{u}$, diffusion $(\nabla\cdot k\nabla)\mathbf{x}$ with diffusion coefficient $k$, and the generation/consumption of a source/sink (reaction) $\mathbf{R}(\mathbf{x},\mathbf{c})$. By explicitly incorporating the spatial gradient $\nabla\mathbf{x}$ of the state field and its divergence $\nabla\cdot\nabla\mathbf{x}$ into the differentiator architecture, we inform PARCv2 with physics priors (Section 3.1). Furthermore, a hybrid integration approach is introduced to leverage the benefits of numerical integration with data-driven integration (Section 3.2), enabling accurate and stable predictions over a long time period. Finally, we also present a two-stage training scheme that separates the process of learning the PDEs and the time integration (Section 3.3).
@@ 3.1 Architecture Design
The dynamics of an advection-diffusion-reaction system can be described by the following form of PDE:
@@ 3.2 Hybrid Integration
For physical systems with unsteady and fast transient characteristics, simulations can quickly deviate from the true dynamic trajectories as numerical and modeling errors accumulate during time integration. Traditional numerical methods for initial value problems such as the Runge-Kutta methods or the Adams methods take multiple steps in time to correct the error accumulation heuristically. However, unsteady and fast transient problems require a higher-order correction with a small time step, making it less tractable for our purpose of rapidly estimating the dynamics. The original PARC Nguyen et al. (2023a) and other works in literature (e.g., Shen et al. (2020)) explored the use of a deep neural network-based quadrature for time integration. However, as we will demonstrate in Sections 4.2 & 4.3, these purely data-driven integrators fail to predict advection features with sharp boundaries precisely.
@@
3.3 Training
@@ 4.1 Burgers’ Equation
In terms of solution quality, measured by the residual of Burgers’ equation from the predicted fields (Table 2), both PARC and PARCv2 demonstrated comparable results to other models. While PhyCRNet and PIFNO exhibited the lowest residuals, likely due to explicit minimization of the PDE residual through physics-informed training loss, the solution quality of PARC and PARCv2 was not significantly inferior, especially considering the ground truth data (DNS) already had a similar PDE residual of 0.1241 cm/s2. The observed high solution quality and low prediction accuracy of PhyCRNet and PIFNO may result from over-regularization due to incorporating the PDE residual term in the training loss. Indeed, this concern has been previously documented in the literature, and the validation results presented here reaffirm its significance.
@@
Finally, it is worthwhile to note that despite choosing a relatively large discrete time step, which may appear to contravene the Courant-Friedrichs-Lewy (CFL) condition of conventional explicit solvers, the stability of the solutions by PARCv2 remains unaffected by this large step size selection. While lacking more solid evidence, we speculate that errors stemming from our discretization choices have been mitigated by both the differentiator and integrator of PARCv2, owing to their robust modeling capabilities.
4.2 Navier-Stokes Equations
@@
Figure 4 displays the fluid velocity fields predicted by the models. Visually, all methods appear to capture the oscillatory vorticity patterns to a certain extent. However, the oscillatory phase of the vorticity patterns was most accurate in PARCv2, highlighting its stable long-time prediction capabilities. This observation is further supported by Table 1, where the RMSE of PARCv2 was 0.0727 m/s—roughly 1/4 of the values obtained by the other methods.
Interestingly, the original PARC, only with a purely data-driven integrator (fourth row in Figure 4), exhibited highly noisy predictions. This may be indicative of the rapid accumulation of network prediction errors in unsteady flows with sharp features. It is also worthwhile to compare PARCv2 (second row in Figure 4) with the original PARC with the numerical integrator (third row in the same figure). Compared to the purely data-driven integration case, the original PARC exhibited a more stable prediction with the numerical integrator. However, the original PARC still appears to be off from the ground truth in terms of the oscillatory phase of the vorticity pattern, while PARCv2 aligns well with the true phase. This result highlights the importance of the data-driven higher-order correction in the integrator. Furthermore, not only the phase but also the geometric pattern slightly deviates from the ground truth in the case of the original PARC, in contrast to the PARCv2 result where the geometric patterns are faithfully reproduced. This is indicative of the role of the spatial gradient terms newly introduced in the PARCv2 architecture.
@@ 4.3 Energy Localization in Energetic Materials
5 Discussion
The enhanced performance found in PARCv2 compared to other baselines across all benchmark problems may be explained theoretically as follows.
Borrowing the views of mathematical physics and differential geometry, an evolving field of a physical quantity can be treated as a trajectory, or a curve, in the Galilean spacetime. A submanifold embedded in the Galilean spacetime would describe some physical rules and constraints. That is, if the physical quantity is subjected to such rules and constraints, the corresponding trajectory or curve must lie on the manifold. Additionally, when there are multiple observations of such evolving fields under different conditions, these observations create a visual of a bundle of curves all of which are lying on the dynamics manifold.
In this setting, PARCv2 can be viewed as learning the topological structures of the dynamics manifold, in contrast to the learning bias approaches that enforce physics via geometric constraints in spacetime. Particularly, PARCv2 attempts to learn the approximation of the tangent spaces on the dynamics manifold (tangent bundle), as well as a “cone” in each tangent space describing where the dynamics may head at a given state. On the contrary, other physics-informed models aim to fit the manifold itself directly, by evaluating the manifold positions across different sample points, or what the literature refers to as “collocation points.”
Given this theoretical view, we argue that the problem of learning the tangent bundle and the tangent cones therein in the case of PARCv2 is a simpler optimization problem, compared to fitting the manifold directly such as in other baselines. The empirical evidence on multiple benchmark problems supports this argument, as discussed in Section 4 about the “tension” between prediction error (RMSE) and solution quality (PDE residuals). When the dynamics manifold is relatively smooth, previous PIML methods may approximate the manifold geometry effectively. However, in the presence of strong nonlinearity and limited data, fitting the solution manifold globally becomes extremely challenging. PARCv2 addresses this by imposing hard constraints on the topology of the dynamics manifold and conducting local physics-informed regularization, allowing for better learning capability in the presence of strong nonlinearity. This approach narrows the search space of PARCv2 learning, resulting in enhanced performance compared to previous PIML methods in handling strong nonlinearity.
@@ 6 Conclusion
These advantages underscore the significance and potential of the inductive bias research direction, positioning it as a promising complement to current mainstream learning bias approaches. It is, in fact, important to note that PARCv2 is not positioned as an exclusive alternative to learning-bias methods, as the incorporation of physics-informed loss during PARC training is still possible. Future research directions may explore the synergistic leverage of soft constraints imposed by the physics-informed loss and the inductive bias provided by the PARC differentiator-integrator architecture. Enhancing the computational efficiency of PARCv2 is also another opportunity for improvement. From our experiments, PARCv2 is slightly heavier than other PIML models, while still significantly less expensive than direct numerical simulations. To this end, potential valuable research directions would include reduced-order modeling (i.e., dimension reduction or learning the latent space dynamics).
@@
Finally, considering the potential societal impacts of PARCv2 is also crucial. For one, interpretability and visualization of PARCv2 is what we deem as the most urgent issue, as it can shed light on how reliable and trustworthy the predictions are. Furthermore, the ability to rapidly explore complex physical phenomena may also be a double-edged sword, as the characterization of sensitive materials, for example, may result in various consequences. Therefore, PARCv2 should be used carefully to avoid any unwanted negative societal impact.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
We appreciate the anonymous reviewers for their comments on improving this work. This work was supported by the Army Research Office (ARO) Energetics Basic Research Center (EBRC) program under Grant No. W911NF-22-2-0164, the National Science Foundation under Grant No. DMREF-2203580, and the Department of Energy under contract No. DE-AC52-07NA27344.
@@ A.1 Burgers’ Equation
Burgers’ equation is a convection-diffusion equation that occurs in a variety of applications such as fluid mechanics, acoustics, plasma physics, and traffic flows. With the presence of the viscous (diffusion) term, the shock discontinuities are smoothed out and one can expect to obtain a well-behaved and smooth solution which will make the learning task less challenging for PARCv2 and other models. Therefore, learning to solve Burgers’ equation can be considered the most straightforward compared to the other two validation studies and was used as the first experiment for our validation.
@@
The fluid parameters, i.e., the Reynolds number $R$ and the initial velocity distribution $a$ and $w$ are varied according to the values listed in Table 3. Note that the fluid parameters in the train-test split are designed to test the models’ abilities to not only interpolate within the training distributions but also extrapolate/generalize to conditions outside the training distributions.
@@
The training data was obtained by solving Eq. 9 with numerical simulation using implicit backward Euler integration. Data was generated using the open-source code provided by Fries et al. [2022]. The original numerical simulation data included 1,500 snapshots and was down-sampled to 100 snapshots to be used for training. The spatial resolution of the numerical simulation data was kept during the training.